OpenStudy (anonymous):

how do u write the standard equation and graph the parabola 12(x+6)=(y-4)^2 ?

OpenStudy (anonymous):

Because y is going to be squared and you want the standard equation ax^2 + bx + c expand (y-4)^2 you should get 12(x+6) = y^2 -8y + 16 now divide both sides by 12 x+6 = (1/12)y^1/2 - (2/3)y + (4/3) now subtract 6 from both sides and your standard equation in terms of y is: x=(1/12)y^1/2 - (2/3)y + (22/3) I sent a really crappy graph of what it will somewhat look like...dont think that is the way it is though, its just an approximation to let you know if yours is right